OpenStudy (anonymous):

A uniform chain is placed on a table with a part of it hanging over the edge. It just begins to slide when it's part is 1/3rd of its length. Find coefficient of friction between table and chain?

OpenStudy (souvik):



OpenStudy (souvik):

let take \(\lambda\) be the mass per unit length.. and l be the total length the the force on 1/3*l is....1/3*l* \(\lambda\)g then f=1/3*l* \(\lambda\)g => \(\mu\)2/3*l* \(\lambda\)g=1/3*l* \(\lambda\)g and \(\mu\)=1/2

OpenStudy (anonymous):

let m be the mass per unit length so hanging part's weight = 1/3 lmg mass of the table chain = 2/3lm F=mewR 1/3lmg=mew* 2/3lmg so mew = 1/2

OpenStudy (anonymous):

got a problem? If not please look at my question too

OpenStudy (anonymous):